If the atomic weight of carbon were set at 24 amu, the value of the Avogadro constant
would be?
1
Expert's answer
2013-01-23T08:02:05-0500
The Avogadro constant enters into the definition ofthe unified atomic mass unit, u,

1 u = Mu / Na = 1,6605 * 10^-24

where Mu is the molar mass constant and Na is Avogadro constant.

Mu = Mw / Ar

If Mw= 24 amu

Mu = 24 amu / 12 = 2

Na' = Mu / 1,6605 * 10^-24 = 2 / 1,6605 * 10^-24 = 1,2
* 10^24

This is " new" Avogadro constant = 1,2 * 10^24